# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v2

@@ 5 Empirical Study: Linear Structure and Understanding the Exploration of PreDeToR
In Figure 2 we first show the linear bandit setting for horizon $n=25$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. Note that the length of the context (the number of rounds) is an artifact of the transformer architecture and computational complexity. This is because the self-attention takes in as input a length-$n$ sequence of tokens of size $d$, and requires $O\left(dn^{2}\right)$ time to compute the output (Keles et al., 2023). Further empirical setting details are stated in Section A.2.
@@
We observe from Figure 2 that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Note that for this low data (short horizon) regime, the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. This results in higher regret. Observe that PreDeToR (-$\tau$) performs quite similarly to LinUCB and lowers regret compared to Thomp which also shows that PreDeToR is able to exploit the latent linear structure and reward correlation of the underlying tasks. Note that LinUCB is close to the optimal algorithm for this linear bandit setting. PreDeToR outperforms AD as the main objective of AD is to match the performance of its demonstrator. In this short horizon, we see that MLinGreedy performs similarly to LinUCB.
@@
We also show how the prediction error of the optimal action by PreDeToR is small compared to LinUCB in the linear bandit setting. In Figure 2(b) we first show how the $10$ actions are distributed in the $M_{\text{test}}=200$ test tasks. In Figure 2(b) for each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then, in Figure 2(c), we show the prediction error of PreDeToR (-$\tau$) for each task $m\in[M_{\text{test}}]$. The prediction error is calculated as $(\widehat{\mu}_{m,n,*}(a)-\mu_{m,*}(a))^{2}$ where $\widehat{\mu}_{m,n,*}(a)=\max_{a}\widehat{\bm{\theta}}_{m,n}^{\top}\mathbf{x}_{m}(a)$ is the empirical mean at the end of round $n$, and $\mu_{*,m}(a)=\max_{a}\bm{\theta}_{m,*}^{\top}\mathbf{x}_{m}(a)$ is the true mean of the optimal action in task $m$. Then we average the prediction error for the action $a\in[A]$ by the number of times the action $a$ is the optimal action in some task $m$. From the Figure 2(c), we see that for actions $\{2,3,5,6,7,10\}$, the prediction error of PreDeToR is either close or smaller than LinUCB. Note that LinUCB estimates the empirical mean directly from the test task, whereas PreDeToR has a strong prior based on the training data. So PreDeToR is able to estimate the reward of the optimal action quite well from the training dataset $\mathcal{D}_{\text{pre}}$. This shows the power of PreDeToR to go beyond the in-context decision-making setting studied in Lee et al. (2023); Lin et al. (2023); Ma et al. (2023); Sinii et al. (2023); Liu et al. (2023c) which require long horizons/trajectories and optimal action during training to learn a near-optimal policy.
@@
We first show in Figure 3(a) the training distribution of the optimal actions. For each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then in Figure 3(b) we show how the sampling distribution of DPT-greedy, PreDeToR  and PreDeToR-$\tau$ change in the first $10$ and last $10$ rounds for all the tasks where action $5$ is optimal. To plot this graph we first sum over the individual pulls of the action taken by each algorithm over the first $10$ and last $10$ rounds. Then we average these counts over all test tasks where action $5$ is optimal. From the figure Figure 3(b) we see that PreDeToR(-$\tau$) consistently pulls the action $5$ more than DPT-greedy. It also explores other optimal actions like $\{2,3,6,7,10\}$ but discards them quickly in favor of the optimal action $5$ in these tasks. This shows that PreDeToR (-$\tau$) only considers the optimal actions seen from the training data. Once sufficient observation have been observed for the task it switches to task-based exploration and samples the optimal action more than DPT-greedy.
Finally, we plot the feasible action set considered by DPT-greedy, PreDeToR, and PreDeToR-$\tau$ in Figure 3(c). To plot this graph again we consider the test tasks where the optimal action is $5$. Then we count the number of distinct actions that are taken from round $t$ up until horizon $n$. Finally we average this over all the considered tasks where the optimal action is $5$. We call this the candidate action set considered by the algorithm. From the Figure 3(c) we see that DPT-greedy explores the least and gets stuck with few actions quickly (by round $10$). Note that the actions DPT-greedy samples are sub-optimal and so it suffers a high cumulative regret (see Figure 2). PreDeToR explore slightly more than DPT-greedy, but PreDeToR-$\tau$ explores the most.
@@ 6 Empirical Study: Importance of Shared Structure and Introducing New Actions
(a) 00 new action
(b) $1$ new action
(c) $5$ new actions
@@
(a) 00 new action
(b) $1$ new action
(c) $5$ new actions
@@
Baselines: We implement the same baselines discussed in Section 4.
@@
Experimental Result: We observe these outcomes in Figure 4 and Figure 5. We consider the linear and non-linear bandit setting of horizon $n=50$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. Here during data collection and during collecting the test data, we randomly select between $0,1,5$, and $10$ new actions from $\mathbb{R}^{d}$ for each task $m$. So the number of invariant actions is $|\mathcal{A}^{\text{inv}}|\in\{10,5,1,0\}$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. From Figure 4(a), 4(b), 4(c), and 4(d), we observe that when the number of invariant actions is less than PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Observe that PreDeToR (-$\tau$) matches LinUCB and has lower regret than DPT-greedy, and AD when $\mathcal{A}^{\text{inv}}|\in\{10,5,1\}$. This shows that PreDeToR (-$\tau$) is able to exploit the latent linear structure of the underlying tasks. However, as the number of invariant actions decreases we see that PreDeToR(-$\tau$) performance drops and becomes similar to the unstructured bandits Thomp. We also show in Section A.3 that in $K$-armed bandit setting when there is no structure across arms PreDeToR (-$\tau$) matches the performance of the demonstrator.
Similarly in Figure 5(a), 5(b), 5(c), and 5(d) we show the performance of PreDeToR in the non-linear bandit setting. Observe that LinUCB, MLinGreedy fails to perform well in this non-linear setting due to their assumption of linear rewards. Again note that PreDeToR (-$\tau$) has lower regret than DPT-greedy, and AD when $\mathcal{A}^{\text{inv}}|\in\{10,1\}$. This shows that PreDeToR (-$\tau$) is able to exploit the latent linear structure of the underlying tasks. However, as the number of invariant actions decreases we see that PreDeToR(-$\tau$) performance drops and becomes similar to AD.
@@ 7 Data Collection Analysis
In this section, we analyze the performance of PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCB when the weak demonstrator $\pi^{w}$ is Thomp, LinUCB, or Uniform. We again consider the linear bandit setting discussed in Section 4. We show the cumulative regret by the above baselines in Figure 6(a), 6(b), and 6(b) when data is collected through Thomp, LinUCB, and Uniform respectively. We first state the main finding below:
@@
Experimental Result: We observe these outcomes in Figure 6. In Figure 6(a) we see that the $A$-actioned Thomp is explorative enough as it does not explore with the knowledge of feature representation. So it pulls the sub-optimal actions sufficiently high number of times before discarding them in favor of the optimal action. Therefore the training data is diverse enough so that PreDeToR (-$\tau$) can predict the reward vectors for actions sufficiently well. Consequently, PreDeToR (-$\tau$) almost matches the LinUCB algorithm. Both DPT-greedy and ADperform poorly in this setting.
In Figure 6(b) we see that the LinUCB algorithm is not explorative enough as it explores with the knowledge of feature representation and quickly discards the sub-optimal actions in favor of the optimal action. Therefore the training data is not diverse enough so that PreDeToR (-$\tau$) is not able to correctly predict the reward vectors for actions. Note that DPT-greedy also performs poorly in this setting when it is not provided with the optimal action information during training. The AD matches the performance of its demonstrator LinUCB because of its training procedure of predicting the next action of the demonstrator.
Finally, in Figure 6(c) we see that the $A$-armed Uniform is fully explorative as it does not intend to minimize regret (as opposed to Thomp) and does not explore with the knowledge of feature representation. Therefore the training data is very diverse which results in PreDeToR (-$\tau$) being able to predict the reward vectors for actions very well. Consequently, PreDeToR (-$\tau$) perfectly matches the LinUCB algorithm. Note that AD performs the worst as it matches the performance of its demonstrator whereas the performance of DPT-greedy suffers due to the lack of information on the optimal action during training.
We also empirically study the test performance of PreDeToR (-$\tau$) in $K$-armed bandit setting when there is no structure across arms in Section A.3, against the original DPT in Section A.3, in other non-linear bandit settings such as bilinear bandits (Section A.4), latent bandits (Section A.5), draw a connection between PreDeToR and Bayesian estimators (Section A.6), and perform sensitivity and ablation studies in Section A.7, A.9, A.10, A.11. Due to space constraints, we refer the interested reader to the relevant section in the appendices.
@@ 8 Theoretical Analysis of Generalization
To show this, we proceed as follows: Suppose we have the training data set $\mathcal{H}_{\mathrm{all}}=\{\mathcal{H}_{m}\}_{m=1}^{M_{\text{pre}}},$ where the task $m\sim\mathcal{T}$ with a distribution $\mathcal{T}$ and the task data $\mathcal{H}_{m}$ is generated from a distribution $\mathcal{D}_{\text{pre}}(\cdot|m).$ For illustration purposes, here we consider the training data distribution $\mathcal{D}_{\text{pre}}(\cdot|m)$ where the actions are sampled following soft-LinUCB (a stochastic variant of LinUCB) (Chu et al., 2011). Given the loss function in Equation 3, we can define the task $m$ training loss of PreDeToR-$\tau$ as ^$\widehat{\mathcal{L}}_{m}(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}})=\frac{1}{n}\sum_{t=1}^{n}\ell(r_{m,t},\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\widehat{r}_{m,t}(I_{m,t})|\mathcal{H}^{t}_{m}))=\frac{1}{n}\sum_{t=1}^{n}(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\widehat{r}_{m,t}(I_{m,t})|\mathcal{H}^{t}_{m})-r_{m,t})^{2}$. We drop the notation $\mathbf{\Theta},\mathbf{r}$ from $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}$ for simplicity and let $M=M_{\text{pre}}$. We define
[…]
@@
(Error stability (Bousquet & Elisseeff, 2002; Li et al., 2023)). Let $\mathcal{H}=\left(I_{t},r_{t}\right)_{t=1}^{n}$ be a sequence in $[A]\times[0,1]$ with $n\geq 1$ and $\mathcal{H}^{\prime}$ be the sequence where the $t^{\prime}$th sample of $\mathcal{H}$ is replaced by $\left(I_{t}^{\prime},r_{t}^{\prime}\right)$. Error stability holds for a distribution $(I,r)\sim\mathcal{D}$ if there exists a $K>0$ such that for any $\mathcal{H},\left(I_{t}^{\prime},r_{t}^{\prime}\right)\in([A]\times[0,1]),t\leq n$, and $\mathrm{TF}\in\mathrm{Alg}$, we have
@@
[…]
@@
(PreDeToR risk) Suppose error stability 8.1 holds and assume loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz for all $r_{t}\in[0,B]$ and horizon $n\geq 1$. Let $\widehat{\mathrm{TF}}$ be the empirical solution of (ERM) and $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ be the covering number of the algorithm space $\mathrm{Alg}$ following Definition C.2 and C.3. Then with probability at least $1-2\delta$, the excess MTL risk of PreDeToR-$\tau$ is bounded by
@@
The proof of Theorem 8.2 is provided in Section C.1. From Theorem 8.2 we see that in low-data regime with a small horizon $n$, as the number of tasks $M$ increases the MTL risk decreases. We further discuss the stability factor $K$ and covering number $\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)$ in Remark C.4, and C.5.
We now present the transfer learning risk of PreDeToR-$\tau$ for an unknown target task $g\sim\mathcal{T}$ with the test dataset $\mathcal{H}_{g}\sim\mathcal{D}_{\text{test}}(\cdot|g)$. Note that the test data distribution $\mathcal{D}_{\text{test}}(\cdot|g)$ is such that the actions are sampled following soft-LinUCB.
Theorem 8.3.
[…]
The proof is given in Section C.2. This shows that for the transfer learning risk of PreDeToR-$\tau$ (once trained on the $M$ source tasks) scales with $\widetilde{O}(1/\sqrt{M})$. This is because the unseen target task $g\sim\mathcal{T}$ induces a distribution shift, which, typically, cannot be mitigated with more samples $n$ per task. A similar observation is provided in Lin et al. (2023). We further discuss this in Remark C.7. We also observe a similar phenomenon empirically; see the discussion in Section A.13.

### v3

@@ 5 Empirical Study: Linear Structure and Understanding PreDeToR’s Exploration
In Figure˜2 we first show the linear bandit setting for horizon $n=25$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. Note that the length of the context (the number of rounds) is an artifact of the transformer architecture and computational complexity. This is because the self-attention takes in as input a length-$n$ sequence of tokens of size $d$, and requires $O\left(dn^{2}\right)$ time to compute the output (Keles et al., 2023). Further empirical setting details are stated in Section˜A.2.
@@
We observe from Figure˜2 that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Note that for this low data (short horizon) regime, the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. This results in higher regret. Observe that PreDeToR (-$\tau$) performs quite similarly to LinUCB and lowers regret compared to Thomp which also shows that PreDeToR is able to exploit the latent linear structure and reward correlation of the underlying tasks. Note that LinUCB is close to the optimal algorithm for this linear bandit setting. PreDeToR outperforms AD as the main objective of AD is to match the performance of its demonstrator. In this short horizon, we see that MLinGreedy performs similarly to LinUCB.
@@
We also show how the prediction error of the optimal action by PreDeToR is small compared to LinUCB in the linear bandit setting. In Figure˜2(b) we first show how the $10$ actions are distributed in the $M_{\text{test}}=200$ test tasks. In Figure˜2(b) for each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then, in Figure˜2(c), we show the prediction error of PreDeToR (-$\tau$) for each task $m\in[M_{\text{test}}]$. The prediction error is calculated as $(\widehat{\mu}_{m,n,*}(a)-\mu_{m,*}(a))^{2}$ where $\widehat{\mu}_{m,n,*}(a)=\max_{a}\widehat{\bm{\theta}}_{m,n}^{\top}\mathbf{x}_{m}(a)$ is the empirical mean at the end of round $n$, and $\mu_{*,m}(a)=\max_{a}\bm{\theta}_{m,*}^{\top}\mathbf{x}_{m}(a)$ is the true mean of the optimal action in task $m$. Then we average the prediction error for the action $a\in[A]$ by the number of times the action $a$ is the optimal action in some task $m$. From the Figure˜2(c), we see that for actions $\{2,3,5,6,7,10\}$, the prediction error of PreDeToR is either close or smaller than LinUCB. Note that LinUCB estimates the empirical mean directly from the test task, whereas PreDeToR has a strong prior based on the training data. So PreDeToR is able to estimate the reward of the optimal action quite well from the training dataset $\mathcal{D}_{\text{pre}}$. This shows the power of PreDeToR to go beyond the in-context decision-making setting studied in Lee et al. (2023); Lin et al. (2023); Ma et al. (2023); Sinii et al. (2023); Liu et al. (2023c) which require long horizons/trajectories and optimal action during training to learn a near-optimal policy.
@@
We first show in Figure˜3(a) the training distribution of the optimal actions. For each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then in Figure˜3(b) we show how the sampling distribution of DPT-greedy, PreDeToR  and PreDeToR-$\tau$ change in the first $10$ and last $10$ rounds for all the tasks where action $5$ is optimal. To plot this graph we first sum over the individual pulls of the action taken by each algorithm over the first $10$ and last $10$ rounds. Then we average these counts over all test tasks where action $5$ is optimal. From the figure Figure˜3(b) we see that PreDeToR(-$\tau$) consistently pulls the action $5$ more than DPT-greedy. It also explores other optimal actions like $\{2,3,6,7,10\}$ but discards them quickly in favor of the optimal action $5$ in these tasks. This shows that PreDeToR (-$\tau$) only considers the optimal actions seen from the training data. Once sufficient observation have been observed for the task it switches to task-based exploration and samples the optimal action more than DPT-greedy.
Finally, we plot the feasible action set considered by DPT-greedy, PreDeToR, and PreDeToR-$\tau$ in Figure˜3(c). To plot this graph again we consider the test tasks where the optimal action is $5$. Then we count the number of distinct actions that are taken from round $t$ up until horizon $n$. Finally we average this over all the considered tasks where the optimal action is $5$. We call this the candidate action set considered by the algorithm. From the Figure˜3(c) we see that DPT-greedy explores the least and gets stuck with few actions quickly (by round $10$). Note that the actions DPT-greedy samples are sub-optimal and so it suffers a high cumulative regret (see Figure˜2). PreDeToR explore slightly more than DPT-greedy, but PreDeToR-$\tau$ explores the most.
@@ 6 Empirical Study: Importance of Shared Structure and Introducing New Actions
(a) 0 new action
(b) 11 new action
(c) 55 new actions
@@
(a) 0 new action
(b) 11 new action
(c) 55 new actions
@@
Baselines: We implement the same baselines discussed in Section˜4.
@@
Experimental Result: We observe these outcomes in Figure˜4 and Figure˜5. We consider the linear and non-linear bandit setting of horizon $n=50$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. Here during data collection and during collecting the test data, we randomly select between $0,1,5$, and $10$ new actions from $\mathbb{R}^{d}$ for each task $m$. So the number of invariant actions is $|\mathcal{A}^{\text{inv}}|\in\{10,5,1,0\}$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. From Figure˜4(a), 4(b), 4(c), and 4(d), we observe that when the number of invariant actions is less than PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Observe that PreDeToR (-$\tau$) matches LinUCB and has lower regret than DPT-greedy, and AD when $\mathcal{A}^{\text{inv}}|\in\{10,5,1\}$. This shows that PreDeToR (-$\tau$) is able to exploit the latent linear structure of the underlying tasks. However, as the number of invariant actions decreases we see that PreDeToR(-$\tau$) performance drops and becomes similar to the unstructured bandits Thomp. We also show in Section˜A.3 that in $K$-armed bandit setting when there is no structure across arms PreDeToR (-$\tau$) matches the performance of the demonstrator.
Similarly in Figure˜5(a), 5(b), 5(c), and 5(d) we show the performance of PreDeToR in the non-linear bandit setting. Observe that LinUCB, MLinGreedy fails to perform well in this non-linear setting due to their assumption of linear rewards. Again note that PreDeToR (-$\tau$) has lower regret than DPT-greedy, and AD when $\mathcal{A}^{\text{inv}}|\in\{10,1\}$. This shows that PreDeToR (-$\tau$) is able to exploit the latent linear structure of the underlying tasks. However, as the number of invariant actions decreases we see that PreDeToR(-$\tau$) performance drops and becomes similar to AD.
@@ 7 Data Collection Analysis
In this section, we analyze the performance of PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCB when the weak demonstrator $\pi^{w}$ is Thomp, LinUCB, or Uniform. We again consider the linear bandit setting discussed in Section˜4. We show the cumulative regret by the above baselines in Figure˜6(a), 6(b), and 6(b) when data is collected through Thomp, LinUCB, and Uniform respectively. We first state the main finding below:
@@
Experimental Result: We observe these outcomes in Figure˜6. In Figure˜6(a) we see that the $A$-actioned Thomp is explorative enough as it does not explore with the knowledge of feature representation. So it pulls the sub-optimal actions sufficiently high number of times before discarding them in favor of the optimal action. Therefore the training data is diverse enough so that PreDeToR (-$\tau$) can predict the reward vectors for actions sufficiently well. Consequently, PreDeToR (-$\tau$) almost matches the LinUCB algorithm. Both DPT-greedy and ADperform poorly in this setting.
In Figure˜6(b) we see that the LinUCB algorithm is not explorative enough as it explores with the knowledge of feature representation and quickly discards the sub-optimal actions in favor of the optimal action. Therefore the training data is not diverse enough so that PreDeToR (-$\tau$) is not able to correctly predict the reward vectors for actions. Note that DPT-greedy also performs poorly in this setting when it is not provided with the optimal action information during training. The AD matches the performance of its demonstrator LinUCB because of its training procedure of predicting the next action of the demonstrator.
Finally, in Figure˜6(c) we see that the $A$-armed Uniform is fully explorative as it does not intend to minimize regret (as opposed to Thomp) and does not explore with the knowledge of feature representation. Therefore the training data is very diverse which results in PreDeToR (-$\tau$) being able to predict the reward vectors for actions very well. Consequently, PreDeToR (-$\tau$) perfectly matches the LinUCB algorithm. Note that AD performs the worst as it matches the performance of its demonstrator whereas the performance of DPT-greedy suffers due to the lack of information on the optimal action during training.
We also empirically study the test performance of PreDeToR (-$\tau$) in $K$-armed bandit setting when there is no structure across arms in Section˜A.3, against the original DPT in Section˜A.3, in other non-linear bandit settings such as bilinear bandits (Section˜A.4), latent bandits (Section˜A.5), draw a connection between PreDeToR and Bayesian estimators (Section˜A.6), and perform sensitivity and ablation studies in Section˜A.7, A.9, A.10, A.11. Due to space constraints, we refer the interested reader to the relevant section in the appendices.
@@ 8 Theoretical Analysis of Generalization
To show this, we proceed as follows: Suppose we have the training data set $\mathcal{H}_{\mathrm{all}}=\{\mathcal{H}_{m}\}_{m=1}^{M_{\text{pre}}},$ where the task $m\sim\mathcal{T}$ with a distribution $\mathcal{T}$ and the task data $\mathcal{H}_{m}$ is generated from a distribution $\mathcal{D}_{\text{pre}}(\cdot|m).$ For illustration purposes, here we consider the training data distribution $\mathcal{D}_{\text{pre}}(\cdot|m)$ where the actions are sampled following soft-LinUCB (a stochastic variant of LinUCB) (Chu et al., 2011). Given the loss function in Equation˜3, we can define the task $m$ training loss of PreDeToR-$\tau$ as $\widehat{\mathcal{L}}_{m}(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}})=\frac{1}{n}\sum_{t=1}^{n}\ell(r_{m,t},\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\widehat{r}_{m,t}(I_{m,t})|\mathcal{H}^{t}_{m}))=\frac{1}{n}\sum_{t=1}^{n}(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\widehat{r}_{m,t}(I_{m,t})|\mathcal{H}^{t}_{m})-r_{m,t})^{2}$. We drop the notation $\mathbf{\Theta},\mathbf{r}$ from $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}$ for simplicity and let $M=M_{\text{pre}}$. We define
$\widehat{\mathrm{TF}}$= $\underset{\mathrm{TF}\in\mathrm{Alg}}{\arg\min}\widehat{\mathcal{L}}_{\mathcal{H}_{\mathrm{all}}}(\mathrm{TF}):=\frac{1}{M}\sum_{m=1}^{M}\widehat{\mathcal{L}}_{m}(\mathrm{TF}),\quad\text{(ERM)}$ (4)
@@
(Error stability (Bousquet & Elisseeff, 2002; Li et al., 2023)). Let $\mathcal{H}=\left(I_{t},r_{t}\right)_{t=1}^{n}$ be a sequence in $A]\times[0,1]$ with $n\geq 1$ and $\mathcal{H}^{\prime}$ be the sequence where the $t^{\prime}$th sample of $\mathcal{H}$ is replaced by $\left(I_{t}^{\prime},r_{t}^{\prime}\right)$. Error stability holds for a distribution $(I,r)\sim\mathcal{D}$ if there exists a $K>0$ such that for any $\mathcal{H},\left(I_{t}^{\prime},r_{t}^{\prime}\right)\in([A]\times[0,1]),t\leq n$, and $\mathrm{TF}\in\mathrm{Alg}$, we have
@@
$\begin{aligned} &\big\lvert\mathbb{E}_{(x,y)}\left[\ell(r,\mathrm{TF}(\widehat{r}(I)|\mathcal{H}))-\ell\left(r,\mathrm{TF}^{\prime}(\widehat{r}(I)|\mathcal{H})\right)-\ell(r,\mathrm{TF}(\widehat{r}(I)|\mathcal{H}^{\prime}))+\ell\left(r,\mathrm{TF}^{\prime}(\widehat{r}(I)|\mathcal{H}^{\prime})\right)\right]\big\rvert\leq\tfrac{K\rho\left(\mathrm{TF},\mathrm{TF}^{\prime}\right)}{n}.\end{aligned}$
@@
(PreDeToR risk) Suppose error stability ˜8.1 holds and assume loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz for all $r_{t}\in[0,B]$ and horizon $n\geq 1$. Let $\widehat{\mathrm{TF}}$ be the empirical solution of (ERM) and $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ be the covering number of the algorithm space $\mathrm{Alg}$ following Definition C.2 and C.3. Then with probability at least $1-2\delta$, the excess MTL risk of PreDeToR-$\tau$ is bounded by
@@
The proof of Theorem 8.2 is provided in Section˜C.1. From Theorem˜8.2 we see that in low-data regime with a small horizon $n$, as the number of tasks $M$ increases the MTL risk decreases. We further discuss the stability factor $K$ and covering number $\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)$ in Remark˜C.4, and C.5. We also present the transfer learning risk of PreDeToR-$\tau$ in Section˜C.2.